Suppose you just bought the new Windoze 8 and
that you realize that it really sucks. As a result you are sure that the price of Microsoft stock (currently trading for $36.4) is going to fall and you tell your broker to short 130 shares.

a. If the broker's initial margin is 50%, what is the value of personal assets you must have in your account to meet the margin requirement?

b. Assume you have the value of personal assets determined in part 1 in your account. If the broker has a maintenance margin of 30%, what is the highest value the price of Microsoft stock can reach before you are issued a margin call?

c. If the price of Microsoft stock falls to $20, what
is the profit you make from the short sale?

2. Suppose you just bought the new Windoze 8 and that you realize that it really sucks. As a result you are sure that the price of Microsoft stock (currently trading for $36.4) is going to fall and you tell your broker to short 130 shares. /a. If the brokers initial margin is 50%, what is the value of personal assets you must have in your account to meet the margin requirement? b. Assume you have the value of personal assets determined in part 1 in your account. If the broker has a maintenance margin of 30%, what is the highest value the price of Microsoft stock can reach before you are issued a margin call? c. If the price of Microsoft stock falls to $20, what is the profit you make from the short sale?
Answer #1

a)total short shares value= 130*36.4=4732
the personal assets we have=50%*4372=2366
b)maintenace margin=30%=(130p-2366)/130p
p=26
margin call will be received if price drops by 26
c)profit we make=final price-initial price
=130*(36.4-20)
=2132